For each ordered pair, determine whether it is a solution to the system of equations. 2x-3y=-8 y=7x+9 Is it a solution? ? (x, for each ordered oair, determine whether it is a solution to the system of equations. clearly state it for all 4 pairs
Answer #1

Solution: Given, 2-3y = -8 y = 7ntg x 3 2-3y = -2 21n-By = -27 subtract 672 + -194 1g NE-1 ya antg = 7(-1) +9 y = 2 Hence, (1